Given: U(x,,x,)-3x 6x2 P 4, P5,1 20 a) Write the function for the indifference curve and graph it. b) Write the function for the budget constraint and graph it c) What are the utility maximizing amounts of xand x2 given the budget constraint:? d) Do the utility maximizing amounts of x and x2 change if P increases to 5 while P stays constant at 5? Why or why not? e) Do the utility maximizing amounts of x and x2 change if P2 increases to 9 while P stays constant at 4? Why or why not?

PLEASE WRITE NEATLY SO I CAN UNDERSTAND AND DIFFERENTIATE THE VARIOUS SYMBOLS AND VARIABLES.
Answer #1

2 XL hirm u (2jK2) = 3x + 6*2 1. Ane Marginal utility of x, i.e. MUI = 20 (X1) Xa) - 3 x =0 and Marginal utility of Xz ie MU2If the consman spmds his entire income on K2, then the untical in tncept of the budget line will be bound by substituting X2=consmption budle is Hunce utility manimizing (x1, 62) - (04) . d Ib Po increases to 5 while P₂ stays at 5 Hace now, mul = 3/8